OpenStudy (unklerhaukus):

\[(1,2)\cap[2,3)=\] @estudier

OpenStudy (unklerhaukus):

can you confirm the solution is the empty set?

OpenStudy (anonymous):

2 included in 1 and not the other so nothing in common.

Parth (parthkohli):

An arbitrary element \(\phi\) will always fulfil the condition \(\phi <2\) in set \((1,2)\). An arbitrary element \(\psi\) will always fulfil the condition \(\psi \ge 2 \) in set \([2,3)\). So, \(\rm \psi \ge 2>\phi\). Therefore, there is no common element.

OpenStudy (anonymous):

They are not really sets....

Parth (parthkohli):

They are. :p

OpenStudy (unklerhaukus):

intervals

Parth (parthkohli):

Intervals are sets.

OpenStudy (unklerhaukus):

an interval in set on an infinite number of points

OpenStudy (anonymous):

@ParthKohli (1,2) What set is this?

Parth (parthkohli):

{1.000000000000.....1, 1.000000......2 ..... 1.99999999999999........9}

OpenStudy (anonymous):

Maybe, it wasn't actually defined...

OpenStudy (unklerhaukus):

\[x\in\mathbb R[1<x<2)]\]

Parth (parthkohli):

The interval (1,2) can be interpreted as a set containing numbers between and not including 1 and 2.

OpenStudy (zarkon):

\[(1,2)=\{x\in\mathbb{R}|1<x<2\}\]

Parth (parthkohli):

Yes, the set builder notation.

OpenStudy (unklerhaukus):

yeah

OpenStudy (unklerhaukus):

thats what i ment @Zarkon

Parth (parthkohli):

In this case, we have \(\phi \) instead.\[(1,2) = \{\phi \in \mathbb R |1<\phi <2\}\]Doesn't make a lot of difference either...

OpenStudy (unklerhaukus):

what is phi?

Parth (parthkohli):

An arbitrary element in \((1,2)\).

Parth (parthkohli):

Or if you're asking it that way, phi is a Greek letter.

OpenStudy (unklerhaukus):

strange choice of dummy variable

Parth (parthkohli):

Greece is a country located in the east, precisely the Asian continent.

OpenStudy (unklerhaukus):

i did not know that

OpenStudy (unklerhaukus):

so i was right? \[(1,2)\cap[2,3)=\emptyset\]

Parth (parthkohli):

It's actually in Eurasia - Europe.

Parth (parthkohli):

Yes, my lad.

OpenStudy (unklerhaukus):

not you

Parth (parthkohli):

I wasn't right? I was.

OpenStudy (anonymous):

I said u were right back on the second post...

OpenStudy (unklerhaukus):

you dont know which way is east parth

Parth (parthkohli):

Eastern Hemisphere. Okay?

OpenStudy (unklerhaukus):

ok

OpenStudy (unklerhaukus):

thanks all,

Parth (parthkohli):

You're welcome.